40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 12 01:33:21 2020, gautam, Update, LSC, Resuming locking activities ExcessFreqNoise_LSC.pdf
    Reply  Fri Mar 13 00:32:38 2020, gautam, Update, LSC, Some progress RFonly_CARM.pngCARMTF.pdf
       Reply  Tue Mar 17 01:22:03 2020, gautam, Update, LSC, Locking updates 
          Reply  Thu Mar 19 03:33:28 2020, gautam, Update, LSC, More locking updates PRFPMI_lock.pngCARMTF.pdfDARM_AvB.pdflockLoss.png
Message ID: 15273     Entry time: Fri Mar 13 00:32:38 2020     In reply to: 15268     Reply to this: 15278
Author: gautam 
Type: Update 
Category: LSC 
Subject: Some progress 

Finally, some RF only CARM, see Attachment #1. During this time, DARM was also on a blend of IR and ALS control, but I couldn't turn the ALS path off in ~4-5 attempts tonight (mostly me pressing a wrong button). Attachment #2 shows the CARM OLTF, with ~2kHz UGF - for now, I didn't bother turning any boosts on. PRCL and MICH are still on 3f signals.

The recycling gain is ~7-8 (so losses >200ppm), but there may be some offset in some loop. I'll look at REFL DC tomorrow.

Can we please make an effort to keep the IFO in this state for the next week or two
- it really helped tonight I didn't have to spend 2 hours fixing some random stuff and could focus on the task at hand.

Attachment 1: RFonly_CARM.png  251 kB  Uploaded Fri Mar 13 01:32:52 2020  | Hide | Hide all
RFonly_CARM.png
Attachment 2: CARMTF.pdf  92 kB  Uploaded Fri Mar 13 01:40:49 2020  | Hide | Hide all
CARMTF.pdf
ELOG V3.1.3-